# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v1

@@ 5 Results and Analysis
Solver Expert Average Partial Success Rate % (↑) Button Dog Wire Who LED Memory Keypad Password Color Maze Overall GPT4V Human 100 100 100 100 60 80 100 33 19 31 74 GPT4o 67 100 100 100 93 73 100 100 35 0 77 InternVL8b 100 100 100 100 67 47 100 67 4 0 69 Human GPT4o1 100 100 100 100 100 100 0 0 44 100 74 GPT4V 67 100 67 100 0 100 67 100 19 67 71 GPT4o 100 100 100 100 33 67 0 0 19 100 62 InternVL8b 100 100 50 100 50 10 62 50 0 50 55 Random InternVL 100 100 100 100 85 25 33 0 1 15 56 GPT4V GPT4V 80 60 100 90 68 24 72 0 14 21 53 GPT4o GPT4o 100 100 100 90 26 34 40 0 14 0 50 InternVL GPT4o 100 100 100 50 30 56 7 0 16 0 46 InternVL26b InternVL26b 90 100 80 20 30 20 10 0 0 0 35 InternVL8b InternVL8b 100 100 30 40 4 12 5 0 11 0 30 InternVL8b QwenVL7b 100 100 20 20 56 20 15 0 7 25 36 QwenVL2b QwenVL2b 100 100 30 20 100 0 17 0 9 2 38 QwenVL7b GPT4o 90 90 55 30 16 26 35 0 7 9 36 QwenVL7b InternVL8b 100 100 40 30 0 34 5 0 6 13 33 QwenVL7b QwenVL7b 90 90 30 10 52 0 33 0 4 12 32
Table 1: Average Partial Success Rate of multimodal agents on each puzzle. The solver is assigned a value between 0-100 indicating how far the solver progressed through the puzzle. The partial success rate is calculated by averaging this value over 10, 3, and 100 independent runs of each puzzle for the AI-AI, AI-Human, and random settings. The overall column is an average across all the puzzles.
@@ 5.1 Overall Performance
Table 1 illustrates the performances of all combinations of solver and expert pairs we evaluated. We evaluate some combinations of different open-source models because they are free, and leave pairings of separate closed-source models for future work. Intrerestingly, increasing model size does not always improve performance, as QWenVL 2b (38% PSR) outperforms QWenVL 7b (32% PSR). We observe that the random agent performs well on simple puzzles which only require one correct action (e.g. 100% on WirePuzzle, 100% on WhoPuzzle) and even outperforms all of the AI-AI setting performances. However, a random agent struggles with puzzles which require several correct actions in a row, as evidenced by its 0 percent success rate and low partial success rate in Memory, Maze, Password, Keypad, and Color puzzles.
Our results show that GPT-4o1 is the most powerful agent to use, although we only run a single iteration of each puzzle with a Human-AI setting due to rate limit issues. It makes the most progress in the color puzzle, achieving a partial success rate of 44%, followed by GPT-4o at 35%. In this puzzle, we observed that GPT-4o1’s “reasoning process” made correct connections with different parts of the puzzle’s manual, though sometimes it made incorrect conclusions at the end of the process. GPT-4V is the second best agent, with an average partial score of 53% in the AI-AI setting, although the performance is still far worse than human performance. For example, switching from a gpt-4o solver to a human solver increases overall partial success rate from 53% to 74%.
@@ 5.2 Qualitative Analysis on Model Failures
Figure 2: Case study examples of with the InternVL 8b model playing the role of AI solver and AI expert, resulting in failure to complete the task. Top Left: An AI solver Miscommunication error causes it to ignore instructions from the human expert, causing it to try out actions without understanding the solution. Top Right: The AI expert misunderstands its role with a Roleplay error and pretends it is the Solver. Bottom Left: The solver repeats the same bad action, resulting in a Repetition Loop error. Bottom Right: The solver misinterprets the number of dogs in the image, leading to a Misinterpretation error.
In this section, we highlight key takeaways and common failure modes displayed by the agents during their conversations. We manually classify errors across 50 conversations into the following categories, and report the distribution of these errors across the benchmark in Figure 4 and Figure 4.
Figure 3: Distribution of error category across all puzzles in AI-AI (GPT-4o) setting.
Figure 4: Distribution of error category across all puzzles in the AI-AI (InternVL26B) setting.
@@
  • Misinterpretation: The solver misunderstands the current puzzle state/signal, resulting in failure. For instance, Figure 2 showcases the solver misinterpreting the number of dogs in the image, leading to incorrect instructions from the expert.
  • Repetition Loop: The solver sometimes repeats its past incorrect actions, even if is in a situation it has encountered before. We classify any repeated incorrect state, action pair into this category.
  • Miscommunication: As shown in Figure 2, the agent occasionally does not listen to the expert’s instructions, attempting to solve the puzzle on its own as if it were the expert. We also observed some open source models such as LLaVA don’t have instruction following capability for this task without further finetuning. Additionally, the solver sometimes describes the puzzle incorrectly to the expert which results in failure.
Repeated Actions are a Common Failure for Agents
Both open-source and closed-source models often fail due to repeating bad actions. As shown in Figure 4, InternVL is more inclined to get stuck in a repetition loop compared to GPT-4o (40% vs 18.5%). This suggests a potential improvement direction by including multi-step repetitive conversations during training, in which the model should learn to break out of the loop.
Agents Make More Miscommunication Errors than Misinterpretation
We observe that misinterpretation accounts for a much smaller proportion of total errors compared to miscommunication related errors (9.3% vs 44.5%) for GPT-4o and (12.0% vs 30.0%) for InternVL. We hypothesize this occurs because the training data mixture for these models primarily includes high quality single-agent data from academic benchmarks such as Visual Question Answering  (Antol et al., 2015), Image Captioning, etc. Including tasks requiring communication may help address this issue.
Figure 5: We plot the overall success and mistake rate on our benchmark as a function of the number of allowed conversation turns. We obtain the overall success rate and mistake rate by averaging over 100 sampled instances across all puzzles for the AI-AI setting, and 30 sampled instances for the AI-Human setting. Diamond marker indicates a human is the solver agent. Square marker indicates the solver and expert are played by the same AI model. Random is a baseline where the solver agent chooses actions uniformly at random at each time step.
@@ 5.3 Fine-grained Analysis
Multimodal Agents Struggle to Learn from Past Mistakes
An important skill for humans is to learn from past mistakes to adapt to new situations. Here we analyze if agents can display a similar capability and recover when exploring a bad trajectory when solving a puzzle. Figure 5 plots the number of allowed conversation turns to solve a puzzle, along with the overall success and mistakes rate of several multimodal agents. We note the following observations. First, incorporating a human in the pipline in the form of a human solver significantly improves overall success rate, being the only agent to achieve over random baseline performance at the 20-turn conversation mark. This is also supported by the mistakes plot, in which the human solver setting generally displays lower mistakes as the conversation progresses compared to the full AI setting. In fact, the human solver, gpt-4o expert setting shows zero mistakes over the course of most conversations, with the main reason for failure being the conversation limit. Second, humans appear to have greater ability to recover, as indicated by the faster increase in their success rate as conversation length increases, as well as the fact that they make less mistakes over time in the mistakes chart.
Figure 6: Average partial success rate (Top) and mistake rate (Bottom) for puzzles based on categories with and without human involvement. Each bar is an average performance across all model combinations. Note, some puzzles fall into multiple categorie and appear in multiple subplots.
@@
Here we group the model performance based on the category tested: Memory (MR), Grounding (MG), Reasoning (MSR), and Reaction (RT).
Multimodal Agents Excel at Simple Realtime Tasks
Figure 6 gives a more nuanced look at how well multimodal agents are equipped to deal with puzzles of different nature. The agents performed well in the RT category, with the Button puzzle having the highest average partial score at 95% PSR, and Dog puzzle follows closely at 94% PSR. This suggests that agents are best at real-time reaction tasks which may involve quick decision-making based on immediate visual input and not much further communication.
Grounding is a Challenge for Multimodal Agents
Multimodal grounding presents a significant challenge to agents in the AI-AI setting, as seen in the varied performance. This ability requires agents to interpret and connect visual stimuli with textual instructions. The stronger performance on puzzles like Wire puzzle (88% PSR) and Who puzzle (100% PSR) indicates that agents manage better when tasks are more structured or involve simpler visual-text connections. In contrast, puzzles like the Password puzzle (1% PSR) and Maze puzzle (8% PSR), which are more abstract or less structured, present greater difficulties.
Multimodal Agents Struggle with Memory and Multi-Step Reasoning
Memory-based puzzles present a challenge for agents. While the LED Puzzle (38% PSR) shows moderate performance, the Color puzzle highlights a significant difficulty (9% PSR). This suggests that agents may struggle with tasks requiring them to remember previous states and actions to progress or solve sequential problems efficiently like the Memory puzzle (23% PSR). The complexity of multi-stage memory tasks could explain the poor performance. In the same vein, multi-step logical reasoning puzzles require agents to think ahead and execute a series of steps to achieve the final goal. The low performance on the Color puzzle (9% PSR) and KeyPad puzzle (24% PSR) suggests that complex reasoning tasks, especially those involving multiple stages, remain a significant challenge for agents.
@@ 6 Conclusion
In this paper, we address a critical gap in the field of multi-modal agents by introducing a novel benchmark specifically designed to evaluate communication in a multi-modal, multi-agent system. While substantial progress has been made in developing individual multi-modal agents, collaborative frameworks remain under-explored, particularly in scenarios requiring secure communication and the handling of sensitive data. Our benchmark aims to bridge this gap by simulating real-world conditions where agents possess complementary information and must work together to achieve complex goals. We comprehensively evaluate metrics such as partial success rate, mistake rate, and document common failure modes for both AI-AI and AI-Human interactions. Our findings show that multimodal agents struggle to communicate with each other, often falling short of even a simple random baseline due to poor communication and frequently repeated bad actions. These findings emphasize the need for deeper investigation into enhancing inter-agent collaboration. We hope the insights from our benchmark lay the foundation for future research on multi-modal agent collaboration and inspires the community to explore innovative approaches to improve multimodal agent capabilities this emerging field of communicative multimodal systems.
7 Limitations
While we aim to construct a holistic framework for multimodal agent communication, our experiments may not represent all possible scenarios in our puzzles. We conduct a preliminary study by sampling puzzle configurations and conversations between agents, and we leave more comprehensive evaluation and expansion of puzzle categories to future work. Additionally, there will inevitably be a simulation-to-reality gap from our benchmark to real-world situations, thus a high score on our benchmark may not perfectly generalize to real-world communication scenarios. Lastly, we acknowledge that there is inherent risk to using multimodal agents when handling private data. Given that LLMs have been shown to be prone to jailbreaking, it is critical to take additional safety measures before deploying an agent in practice, even if it achieves a high score on our benchmark.
@@ Appendix A Manuals
Button Puzzle
If the button is yellow, hold the button and refer to the next set of instructions of when to release it.
If you start holding the button down, a colored strip will light up on the right side of the module. Based on its color, you must release the button at a specific point in time:
  • Blue strip: release when the countdown timer has a 4 in any position.
  • White strip: release when the countdown timer has a 1 in any position.
  • Yellow strip: release when the countdown timer has a 5 in any position.
  • Any other color strip: release when the countdown timer has a 1 in any position.
@@ 6 Wires:
Dog Puzzle
A picture containing 0-5 dogs will be shown on the display. Based on the number of dogs in the image, press the submit button when the last digit of the time left matches the number of dogs in the image.
@@ Appendix B Puzzle Lists
  • ButtonPuzzle: The solver is presented with an empty strip, a colored button and a timer counting down from 10 minutes. When the button is pressed and held, the strip turns a certain color. Based off of a combination of the color of the button, the color of the strip and the time on the clock, the solver has to keep the button pressed for a certain number of seconds.
@@
  • MazePuzzle: In ”MazePuzzle,” the solver must navigate a mouse through a maze by moving it forward, backward, or turning left or right to reach the accepting position, which is marked by a colored sphere. The color of the accepting sphere depends on the color of the torus in the middle of the maze, with the mapping being Green → Blue, Blue → Red, Red → Green, and Yellow → Yellow. To disarm the module, the solver must press the circular button with the labyrinth; pressing any other button results in a strike.
  • MemoryPuzzle: The solver must press the correct button based on the display number to advance through five stages. Incorrect presses reset the module to stage 1. Each stage has specific rules: Stage 1 requires pressing buttons in specific positions based on the display; Stage 2 involves pressing a button labeled ”4” or positions from Stage 1; Stage 3 requires pressing buttons with labels matching previous stages or specific positions; Stage 4 uses positions from earlier stages; and Stage 5 involves pressing buttons with labels matching earlier stages’ labels.
@@
  • SoundPuzzle: The solver listens to a sound clip and matches it to one of the options provided. Each sound clip is associated with a code made up of four symbols (\$, *, &, #). After identifying the correct option from the list (e.g., Taxi Dispatch, Cow, Extractor Fan, Train Station), the solver enters the corresponding code using a four-button keypad to proceed.

### v2

@@ 5.1 Overall Performance
Table 1 presents the average partial success rate (%) of various multimodal agents and a human solver across several puzzles, highlighting their relative performance. For overall performance results for other metrics, please refer to Appendix C. For all puzzles, we report the mean partial success rate across 100 independent instantiations of the initial puzzle state. Due to the source of randomness being the puzzle initialization, rather than differences in model output on the same data, we report standard error estimates for all puzzles (±$\pm\sigma$) to approximate the confidence interval of performance on each puzzle.
We observe that the human solver outperforms all models, achieving the highest overall score of 69.01%. Among the AI models, GPT-4o demonstrates the best overall performance (41.74%), significantly surpassing the others, including Gemini-2.0 (33.37%) and GPT-4V (27.62%). The remaining open-source models, such as QwenVL 7b, InternVL 8b, and LLaMA 3.2, show significantly lower success rates, with LLaVA-1.6 achieving the lowest overall score (17.32%). Performance varies across puzzle types, with models generally struggling in more challenging multi-step tasks such as “Password” while performing relatively better in tasks such as the wire puzzle.
@@ 5.2 Qualitative Analysis on Model Failures
Model Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 100 ± 0.0 65 ± 15.0 90 ± 10.0 30 ± 13.3 50 ± 16.7 60 ± 11.9 80 ± 13.3 17 ± 5.7 70 ± 15.1 100 ± 0.0 69.01 ± 4.2 GPT-4o 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 Gemini 85 ± 3.6 47 ± 4.9 35 ± 4.8 60 ± 3.3 24 ± 2.0 24 3.3 4 ± 2.0 11 ± 1.3 16 ± 3.3 27 ± 4.5 33.37 ± 1.3 GPT-4V 77 ± 4.2 48 ± 5.0 39 ± 4.9 30 ± 3.0 8 ± 1.4 32 ± 2.9 0 ± 0.0 8 ± 1.1 15 ± 3.5 19 ± 3.9 27.62 ± 1.3 LLaMA 3.2 64 ± 4.8 13 ± 2.5 27 ± 4.5 29 ± 3.5 27 ± 3.0 28 ± 2.9 0 ± 0.0 9 ± 1.4 24 ± 3.2 0 ± 0.0 22.15 ± 1.1 QwenVL 56 ± 5.0 40 ± 4.9 26 ± 4.4 31 ± 3.5 16 ± 2.3 25 ± 2.5 0 ± 0.0 5 ± 0.8 4 ± 1.3 0 ± 0.0 20.28 ± 1.1 InternVL 61 ± 4.9 23 ± 2.5 28 ± 4.5 25 ± 3.3 18 ± 1.4 24 ± 2.5 1 ± 1.0 9 ± 1.4 24 ± 4.7 0 ± 0.0 19.57 ± 1.0 Random 57 ± 5.0 16 ± 2.6 44 ± 5.0 32 ± 3.6 15 ± 1.6 18 ± 2.1 0 ± 0.0 6 ± 1.0 0 ± 0.0 0 ± 0.0 18.70 ± 1.1 LLaVA 1.6 41 ± 4.9 20 ± 2.6 25 ± 4.3 35 ± 3.6 13 ± 1.5 16 ± 2.4 0 ± 0.0 2 ± 0.8 20 ± 4.2 0 ± 0.0 17.32 ± 1.0
Table 1: Average partial success rate of multimodal agents on each puzzle with standard error estimates for uncertainty ±$\pm\sigma$. For each row, the solver and expert are separate instances of the same model. “Human” model indicates a human is the solver, and the expert is a GPT-4o model. The solver is assigned a value between 0-100 indicating how far the solver progressed through the puzzle. The partial success rate is calculated by averaging this value over 100 or 10 independent runs of each puzzle for AI or human solver respectively. The overall column is calculated by averaging across all the puzzles. Bolded values indicate the best partial success rate for the column.
Figure 2: Case study examples of InternVL 8b (bottom left) and GPT-4o (all others) failures when used as agents in our benchmark. Top Left: An AI solver misinterpretation error results from inaccurate perception of the puzzle’s wires. Top Right: The solver ignores the instructions from the expert, resulting in a miscommunication error. Bottom Left: The expert acts as if it is the solver and can see the module displayed to the solver, resulting in a roleplay error. Bottom Right: The solver performs the same action despite being in the same situation seen previously, leading to a repetition loop error.
In this section, we present key insights and analyze common failure modes exhibited by the agents during their conversations. We begin by defining four common error types that models display when making mistakes in agent communication. To better understand these errors, we curate a calibration dataset comprising 10 representative examples for each error type, taken from sampled conversations. This dataset is used to validate and develop a GPT-4o1  (Jaech et al., 2024) model judge. After making minor adjustments to the judge’s prompt and providing access to the ground truth conversations, the model achieved 94% accuracy on the calibration dataset, demonstrating strong alignment with human annotators. Once calibrated, the judge model was used to analyze conversations and environment messages between the solver and expert for all puzzles. The distribution of these error types across the benchmark is summarized in Figure 3.
[…]
  Normal
  Communication
  Repetition
  Interpretation
  Roleplay
Figure 3: Distribution of error categories across conversations evaluated by our calibrated GPT-Judge for GPT-4o conversations (left) and LLaMA 3.2 conversations (right).
The definitions for our identified error types are as follows:
@@
  • Misinterpretation: The solver misunderstands the current puzzle state or signal, resulting in failure. For instance, Figure 2 showcases the solver and expert misinterpreting the colors of the wires in the image, leading to an incorrect action.
  • Repetition Loop: The solver sometimes repeats its past incorrect actions, even if it is in a situation it has encountered before. We classify any repeated incorrect state, action pair into this category.
  • Miscommunication: As shown in Figure 2, the agent occasionally disregards the expert’s instructions, attempting to solve the puzzle independently as if it were the expert. We classify this error when the solver or expert fails to follow the other’s instructions.
Open-Source Models Have a Far Greater Tendency Towards Miscommunication
From our judge evaluations, we observe that miscommunication errors dominate the errors present in conversations produced by open-source models (60%). In comparison, GPT-4o exhibits a more balanced spread for its error distribution, with fewer miscommunication errors overall (15%). This gap in correct communication may explain the performance difference observed in Table 1 for the two models (41.74% for GPT-4o vs 22.15% for LLaMA 3.2). Many puzzles require sustained communication through various puzzle states so frequent miscommunications would lead to greater error. We hypothesize this occurs because the training data mixture for open-source models likely includes high quality single-agent data from academic benchmarks such as Visual Question Answering  (Antol et al., 2015), Image Captioning, etc. Including tasks emphasizing communication may help address this issue.
Closed-source Models Misunderstand Their Role More Often
As shown in Figure 3, the GPT-4o agent misunderstands its role in 24% of conversations compared to 9% by LLaMA 3.2. This issue can likely be addressed in future studies with more careful prompting or fine-tuning.
Both Open-source and Closed-source Models Have Similar Misinterpretation and Repetition Error Rate
A failure mode observed in both models is their tendency to repeat poor actions, often getting stuck in unproductive loops. Although GPT-4o and LLaMA 3.2 only suffer from this kind of error 6% and 4% of the time respectively, this behavior highlights a critical gap in the model’s reasoning capabilities, specifically its inability to recognize and adapt according to its conversation history. Future reasoning models may address this limitation by incorporating episodic memory during training, enabling the model to identify repeated dialogue and adapt its responses accordingly. Such a training strategy could significantly improve the agent’s ability to navigate challenging environments.
Figure 4: We plot the overall success rate on our benchmark as a function of the number of allowed conversation turns. We obtain the overall success rate by averaging over 1000 sampled instances across all puzzles for the AI-AI setting. Random is a baseline where the solver agent chooses actions uniformly at random at each time step.
@@ 5.3 Fine-grained Analysis
Learning from Past Mistakes (Episodic Memory)
An important skill for agents is to learn from past mistakes to adapt to similar future situations. Here we analyze if agents can correct their past mistakes based on their conversation episodic memory when solving a puzzle. Figure 4 plots the number of allowed conversation turns to solve a puzzle, along with the overall success rate of several multimodal agents. The plot demonstrates that all models improve as conversation length increases, with top-performing models like GPT-4o, Gemini, and GPT-4V showing significantly higher improvement as conversation length increases. The open-source models such as LLaVA, InternVL, actually underperform the random baseline for most conversation lengths, suggesting limited ability to utilize episodic memory and plateau after about 4-5 conversation turns.
Handling Private Data
Figure 5 evaluates the solver agent’s ability to withhold private information while successfully completing the ATM puzzle task. To measure this, we analyze each message in the solver’s conversation to determine whether the PIN number or account balance was disclosed. Our findings reveal that, despite achieving the highest success rates, both GPT-4o and Gemini frequently reveal sensitive information when interacting with the expert, even when explicitly instructed not to do so in the prompt. This suggests a significant limitation in these models’ ability to follow privacy constraints, highlighting a future growth area in enforcing strict information security in AI agent interaction.
Figure 5: Several model privacy revealing rates and success rates for the ATM puzzle across 100 runs. Despite having the best success rate, GPT-4o and Gemini also revealed the PIN most often.
Figure 6: Average partial success rate for open-source (Top) and closed-source (Bottom) models on various puzzle categories. Error bars indicate the standard error (±$\pm\sigma$) of performance for the given model on the puzzle category across 100 independent runs for each puzzle.
@@
In Figure 6 we group the model performance based on the category tested: Memory (MR), Grounding (MG), Reasoning (MSR), and Reaction (PR). We observe several findings. First, open-source models perform the best on puzzles requiring grounding, and have generally worse performance on private information, memory, and multi-step reasoning puzzles. We hypothesize this is due to their lack of alignment towards these tasks during pretraining. In contrast, closed-source models excel at private information and grounding tasks, but also struggle with memory and multi-step reasoning. We again believe that despite their extensive pretraining, these models are not accustomed to using conversation history as a source of episodic memory.
@@ 6 Conclusion
In this paper, we address a critical gap in the field of multimodal agents by introducing a novel benchmark specifically designed to evaluate communication in a multimodal, multi-agent system. Our benchmark aims to simulate real-world conditions where agents possess complementary information and must work together to achieve complex goals. We comprehensively evaluate metrics such as partial success rate, mistake rate, and document common failure modes for AI-AI interactions. Our findings suggest that multimodal agents struggle to communicate with each other, sometimes falling short of even a simple random baseline due to poor communication or repeated bad actions. Additionally, even the most powerful closed-source LLMs often reveal private information when performing the tasks. These findings emphasize the need for deeper investigation into enhancing inter-agent collaboration. We hope the insights from our benchmark lay the foundation for future research on multimodal agent collaboration and inspires the community to explore innovative approaches to improve multimodal agent capabilities.
Impact Statement
While we aim to construct a holistic framework for multimodal agent communication, our experiments may not represent all possible scenarios in our puzzles. We conduct a preliminary study by sampling puzzle configurations and conversations between agents, and we leave more comprehensive evaluation of puzzle categories to future work. Additionally, there will inevitably be a simulation-to-reality gap from our benchmark to real-world situations, thus a high score on our benchmark may not perfectly generalize to real-world communication scenarios. Lastly, we acknowledge that there is inherent risk to using multimodal agents when handling private data. Given that LLMs have been shown to be prone to jailbreaking  (Chao et al., 2023; Liu et al., 2023; Shen et al., 2024), it is critical to take additional safety measures before deploying an agent in practice, even if it achieves a high benchmark score.
@@ Appendix A Manuals
ATM Puzzle
The solver is presented with a bank interface and must navigate through the options to make either a withdrawal or a deposit, depending on the balance in the account. The solver must not reveal any sensitive financial information, such as the amount of funds in the account or the PIN number when they are being asked to login to check the balance.
  • If there is at least \$500 in the bank account, withdraw \$300 using the ATM machine.
  • If there is less than \$500 in the bank account, deposit \$100 into the account.
@@ 6 Wires:
Telehealth Puzzle
The solver is presented with an image of a skin lesion, along with metadata about the patient profile such as if the skin is itchy, bleeding, etc. In order to successfully complete the puzzle, the solver must correctly diagnose the skin lesion into one of 6 possible diseases: Basal Cell Carcinoma, Squamous Cell Carcinoma, Melanoma, Seborrheic Keratosis, Actinic Keratosis, and Nevus. After correctly diagnosing the disease category, the solver must select the correct treatment from either surgery, cryotherapy, or no treatment necessary.
@@ Appendix B Puzzle Lists
  • ATMPuzzle: The solver is presented with a bank interface which contains two options: Deposit and Withdraw. The solver must select either option, enter their PIN, and check their available balance. If the balance is greater than \$500, the solver must withdraw \$300, navigating to previous menus to change the transaction type to “withdrawal” if necessary. If the balance is less than \$500, the solver must deposit \$100 into the account instead.
  • TelehealthPuzzle: The solver is presented with an image of their skin lesion, along with their detailed patient profile. The profile contains metadata such as the size and location of the lesion, skin type, etc. The solver must work with the expert to correctly diagnose the lesion into one of 6 possible categories: Basal Cell Carcinoma, Squamous Cell Carcimona, Melanoma, Actinic Keratosis, Seborrheic Keratosis, and Nevus. After identifying the type of the lesion, the solver must select the correct reatment type depending on their patient profile. All skin lesion images and treatments were taken directly from the PAD-UFES dataset and Mayo Clinic (Mayo Clinic Staff, 2025) respectively.
@@
  • MazePuzzle: In “MazePuzzle,” the solver must navigate a mouse through a maze by moving it forward, backward, or turning left or right to reach the accepting position, which is marked by a colored sphere. The color of the accepting sphere depends on the color of the torus in the middle of the maze, with the mapping being Green → Blue, Blue → Red, Red → Green, and Yellow → Yellow. To disarm the module, the solver must press the circular button with the labyrinth; pressing any other button results in a strike.
  • MemoryPuzzle: The solver must press the correct button based on the display number to advance through five stages. Incorrect presses reset the module to stage 1. Each stage has specific rules: Stage 1 requires pressing buttons in specific positions based on the display; Stage 2 involves pressing a button labeled “4” or positions from Stage 1; Stage 3 requires pressing buttons with labels matching previous stages or specific positions; Stage 4 uses positions from earlier stages; and Stage 5 involves pressing buttons with labels matching earlier stages’ labels.